What is the percent yield if 14.0 grams of NaC𝑙(𝑠) is recovered from the reaction of 9.16 grams of C𝑙₂(𝑔) according to the reaction below?


2Na(s)+Cl2(g)β€”>2NaCl(s)


1
Expert's answer
2021-04-19T05:07:27-0400

From the balanced equation

71g of Cl2 yields 2(58.5)g of NaCl

9.16g of Cl2 will yield

117.0/71 x 9.16= 15.09g


% yield = 9.16/15.09 x 100% = 60.7%